# Probing Self-Interacting Dark Matter via Gravitational-Wave Background from Eccentric Supermassive Black Hole Mergers

### v1

@@ Abstract
The nature of dark matter is still mysterious despite various astronomical evidence. As a possible candidate, self-interacting dark matter (SIDM) can potentially resolve some issues appearing in cold dark matter paradigm. Here we investigate how SIDM around supermassive black holes (SMBH) in galaxy centers may form a density spike and imprint in the spectrum shape of stochastic gravitational-wave background from SMBH binaries (SMBHBs). Employing a refined dynamical friction formula and consistently evolving the orbital dynamics, we demonstrate that current pulsar timing arrays (PTAs) data is sensitive to the cross section of SIDM with $\sigma(v)/m_{\chi}\lesssim 0.66\,\mathrm{cm}^{2}/\mathrm{g}$, comparable to other astrophysical probes. We also highlight the importance of including the eccentricity of SMBHBs in the parameter inference, which would affect the results significantly. Our findings reveal the promising potential of PTAs observations in probing the nature of dark matter.
@@ I Introduction
Dark matter is one of the main components of the Universe, with its existence firmly established through multiple astronomical observations including galaxy rotation curves [1, 2], Bullet Cluster [3, 4], cosmic microwave background (CMB) anisotropies [5], and large-scale structure [6]. Although at galactic scale and beyond these observations are consistent with the cold dark matter (CDM) paradigm, at small scales or near galactic centers, the particle nature dark matter remain uncertain. Among numerous dark matter candidates, self-interacting dark matter (SIDM) is a popular and attractive model that may behave differently at small scales [7, 8, 9, 10]. While no definitive signatures have been detected in terrestrial experiments, any new proposals to probe the nature of dark matter would be warranted.
Since the first detection of gravitational waves (GWs) [11], the advent of gravitational-wave astronomy has opened new detection windows for dark matter. Studies of GW signals with both ground‑based and space‑based detectors—such as LVK [12, 13, 14], LISA [15], and Taiji [16]—enable discrimination among a variety of dark‑matter candidates, such as CDM [17, 18], SIDM [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], and ultralight dark matter(ULDM) [31, 32, 33, 34, 35, 36, 37]. GWs detection has demonstrated remarkable potential for fundamental physics [38, 39, 40, 41, 42, 26]. Notably, the search for dark matter is expected to become increasingly intertwined with gravitational-wave astronomy.
@@
While the canonical picture predicts that the characteristic strain $h_{c}$ should scale as $f^{-2/3}$ for a population of inspiraling SMBHBs [49, 50], environmental effects around black holes may alter the merger process, causing deviations of the stochastic gravitational-wave background from the ideal power-law spectrum. NANOGrav study suggest that environmental effects may play a significant role in supermassive black hole mergers [51]. For instance, surrounding matter components such as gas, stars, and dark matter can accelerate the energy dissipation of binary systems. More studies have investigated the effects of three-body interactions [52], CDM [53, 54], ULDM [55, 56, 57, 58] and eccentric orbits on the stochastic gravitational-wave background at low frequencies [59].
@@
In this work, we adopt the SIDM framework with a velocity‑dependent scattering cross section and incorporate a refined dynamical friction formulation that captures additional drag from high‑velocity dark matter particles. We then investigate how dynamical friction affects the eccentric orbits of SMBHBs, which not only augments gravitational radiation but also induces a characteristic suppression in the SGWB’s spectrum at low frequencies relative to circular binaries. Our analysis furnishes a more realistic physical description. Through Bayesian inference applied to observed characteristic‑strain data, we obtain the parameter fit range of SIDM model and quantify how initial orbital eccentricity reshapes the posterior distributions of model parameters.
This paper is organized as follows. Sec. II presents the methodology for constructing dark matter density and velocity profiles around binary SMBHs, and we compare density profiles under different parameter configurations. Sec. III details numerical computations of binary SMBH evolution incorporating eccentric orbits and dynamical friction, with particular emphasis on quantifying contributions from additional drag terms. Sec. IV demonstrates how variations in SIDM parameters and initial orbital eccentricity affect the characteristic strain spectrum of stochastic gravitational waves, comparing these predictions with PTA observations. In Sec. V we employ Bayesian inference and $\chi^{2}$ analysis to quantify parameter compatibility with observational data, ultimately showcase the sensitive parameter space of SIDM models.
@@ II.1 OUTER HALO DENSITY
We consider spherically symmetric dark matter halos and use the Navarro-Frenk-White (NFW) profile [64] to characterize dark matter distribution in the outer region as a function of radial distance $r$ from galaxy center,
@@
[…]
@@
where $M_{\mathrm{ref}}$ is defined as $10^{12}\,h^{-1}M_{\odot}$, and the remaining parameters $C_{c},\gamma_{c},M_{c}$ are expressed as functions of redshift. We have compiled the polynomials mentioned here, along with those referenced in Eq. (6), in Appendix. A. Once we obtain $r_{s}$, we can inversely solve for $\rho_{s}$ through the Eq. (7).
@@ II.2 DENSITY AND VELOCITY DISTRIBUTION OF SIDM
Here $v_{t}$ denotes the transition velocity, defined as the mass ratio between mediator and dark matter $v_{t}\equiv m_{A^{\prime}}/m_{\chi}$ and $v$ denotes the velocity dispersion of the dark matter particles. Here we does not specify the particle physics model of dark matter [70, 71, 72, 73, 74, 75, 20, 74, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87], as the relevant physical parameters are represented by $\sigma_{0}$ and $v_{t}$. It can be seen that the cross section behaves like a constant at low velocities, while at high velocities it transitions to a behavior similar to Coulomb scattering, $\sigma(v)\propto 1/v^{4}$.
Due to the large cross-section of SIDM, the exchange of momentum and energy among particles within a SIDM halo occurs rapidly. In the region with radial coordinate $r$ smaller than some value $r_{1}$, the dark matter density tends to flatten, forming an isothermal core. The boundary $r_{1}$ of the isothermal core is defined through the average interaction cross-section [19, 62],
@@
where $t_{\mathrm{age}}$ represents the age of galaxy core. This indicates that dark matter particles within the radius $r_{1}$ have at least one interaction during the galaxy merger process. We use $t_{\mathrm{age}}=1\,\mathrm{Gyr}$ in this paper for convenience. This does not affect our final analysis of the parameter $\sigma_{0}/m_{\chi}$, as we can easily rescale to different $t_{\mathrm{age}}$ by adjusting the size of $\sigma_{0}/m_{\chi}$ and compute the cross section for different characteristic velocities.
@@
where $v_{0}$ represents the velocity dispersion in isothermal core.
@@
Figure 1: We take the parameters of the binary supermassive black hole system as example $M_{\mathrm{tot}}=10^{10}\,M_{\odot}$, $z=0$, $q=0.5$. Upper: The dark matter density profile is plotted for different transition velocities while fixing the dark matter self-interacting cross section $\sigma_{0}/m_{\chi}=1\,\mathrm{cm}^{2}/\mathrm{g}$. Lower: The dark matter density profile with different cross section while fixing the transition velocity $v_{t}=500\,\mathrm{km}/\mathrm{s}$. Here, the region $r<10^{2}\,\mathrm{pc}$ corresponds to the spike, the flat middle region is the isothermal core, and the $r>10^{6}\,\mathrm{pc}$ segment belongs to the NFW region.
@@
This also represents the radius of the dark matter spike. For CDM, the density distribution of the dark matter spike can be described by a power law [90, 91, 43]. In the case of SIDM, under the non-relativistic approximation, the density and velocity distribution can also be expressed in the form of a power law [92, 93, 62],
@@
The exponent $a$ represents the dependence of the interaction cross-section on the velocity power. For the dark photon model, the interaction cross-section approximates to a constant $(a=0)$ at low velocities, while at high velocities, $a=4$.
@@
We note that for certain masses of central black holes and redshifts, only a sufficiently large interaction cross-section $\sigma_{0}/m_{\chi}$ can ensure that the radius of the dark matter spike $r_{\mathrm{sp}}$ exceeds the radius $r_{1}$ of the isothermal core. For dark matter spikes that have lost their isothermal core, the stable structure may be disrupted due to perturbations caused by the loss of orbital energy during binary inspiral. Meanwhile, very large cross section may also lead to core collapse or a weakening of dynamical friction effects, causing the binary merger timescale to exceed the Hubble time [62, 94, 61].
@@
We define the transition radius $r_{t}$ such that $v(r_{t})=v_{t}$, the density distribution within the spike is given by
$\rho_{\mathrm{sp}}=\begin{cases}\rho_{0}\left(\frac{r_{\mathrm{sp}}}{r_{t}}\right)^{\frac{3}{4}}\left(\frac{r_{t}}{r}\right)^{\frac{7}{4}},&\text{if }r<r_{t}\\ \rho_{0}\left(\frac{r_{\mathrm{sp}}}{r}\right)^{\frac{3}{4}}.&\text{if }r>r_{t}\end{cases}$ (16)
@@
In Fig. 1, we construct the complete density profile and examine how the density profile responds to varying transition velocities $v_{t}$. A distinct flattened density region, corresponding to the isothermal core, appears between approximately $10^{2}$ and $10^{4}\,\mathrm{pc}$. For $v_{t}=300\,\mathrm{km}\,\mathrm{s}^{-1}$, the one–dimensional velocity dispersion is $v_{0}\simeq 373\,\mathrm{km}\,\mathrm{s}^{-1}$, so $v_{0}>v_{t}$. The corresponding transition radius $r_{t}$ lies outside the spike radius $r_{\mathrm{sp}}$, where the density slope is $\gamma=7/4$. Beyond $\sim 10^{6}\,\mathrm{pc}$, the profile reverts to the NFW form. In the lower panel of Fig. 1, with $v_{t}=500\,\mathrm{km}\,\mathrm{s}^{-1}$ fixed, we can see that increasing the cross section would broaden the flat core, which boosts particle velocities and thus lowers the density.
@@
To verify the consistency between our model and observational results, we also conduct a benchmark analysis using the Milky Way system. By cross-validating our theoretical predictions with observational constraints derived from monitoring the $S2$ stellar orbit around the Galactic Center supermassive black hole $\mathrm{Sgr}\mathrm{A}^{*}$, we find that the enclosed mass within the $S2$ orbital pericenter, even with the parameters $v_{t}=100\,\mathrm{km}/\mathrm{s},\,\sigma_{0}/m_{\chi}=0.3\,\mathrm{cm}^{2}/\mathrm{g}$ that yield the highest spike density, the density at $r=0.01\,\mathrm{pc}$ is approximately $10^{7}\,M_{\odot}/\mathrm{pc}^{3}$, and the enclosed mass does not exceed $100\,M_{\odot}$, one order of magnitude smaller than the empirically derived upper limit of $1200\,M_{\odot}$ [95].
@@ III DYNAMICAL FRICTION AND ORBITAL EVOLUTION
where $M_{\mathrm{bh}}$ represents the black hole mass, $v_{\mathrm{bh}}$ denotes the black hole velocity and $\Lambda$ is the Coulomb logarithm, which we take $\log\Lambda=3$ in this work [98]. $N_{1}$ and $N_{2}$ represent the fraction of dark matter particles with velocities less than the black hole’s velocity and that with velocities between the black hole’s velocity and the escape velocity, respectively.
@@
Figure 2: This figure shows the $e(t)$ versus $a(t)$ during the dynamical friction phase of binary supermassive black holes with parameters $M_{\mathrm{tot}}=10^{10}\,M_{\odot}$, $q=0.6$, $z=2$, $v_{t}=300\,\mathrm{km}/\mathrm{s}$, and $\sigma_{0}/m_{\chi}=6\,\mathrm{cm}^{2}/\mathrm{g}$. The dashed curves represent the evolution according to the classical Chandrasekhar formula ($N_{2}=0$), while the solid curved have included the correction term $N_{2}$. The differently colored curves in the plot correspond to different initial eccentricities.
Treating all orbits as circular would undoubtedly neglect the modifications in gravitational radiation induced by eccentric orbits, thereby leading to an overestimation of the contribution from dynamical friction. In this paper, we consider orbital evolution with non-zero eccentricity, as opposed to the simplified circular orbit scenario.
@@
We use the phase-averaging method to approximate the change rates of $i$-th object’s energy and angular momentum in the binary system due to dynamical friction,
@@
The evolution of the orbital semi-major axis and eccentricity as functions of time are determined by the variation of the total energy and angular momentum. Detailed derivations are provided in Appendix. B.
@@
We select supermassive black hole merger events within a mass range of $M_{\mathrm{tot}}/\,M_{\odot}\in[10^{6},10^{11}]$, assuming that the inspiral process of the eccentric orbit begins at an average orbital frequency $F_{0}=10^{-13}\,\mathrm{Hz}$ [100, 101]. High-resolution cosmological $N$-body simulations (IllustrisTNG100-1) [102] indicate that approximately $95\%$ of galactic merger orbits exhibit high eccentricities [101, 103], with binary systems maintaining eccentricities distributed between 0.2 and 0.8 even as they evolve into the PTA frequency band. Therefore, we consider a binary evolution system with different initial eccentricity distribution. Using numerical methods, we can determine the values of the semi-major axis $a(t)$ and eccentricity $e(t)$ at any subsequent time.
@@
Figure 3: Left: We present the variation of characteristic strain with dark matter model parameters using a single source as an example. The source parameters are $M_{\mathrm{tot},0}=10^{10}\,M_{\odot}$, $z_{0}=2$, $q_{0}=0.6$, with the event rate given by $\mathrm{d}^{3}N/(\mathrm{d}M_{\mathrm{tot}}\mathrm{d}z\mathrm{d}q)=\delta(M_{\mathrm{tot}}-M_{\mathrm{tot},0})\delta(z-z_{0})\delta(q-q_{0})$. In the figure, curves of different colors represent characteristic strain with different transition velocities and cross sections, where dashed ones correspond to an initial eccentricity of $0.8$, while the solid represent an initial eccentricity of $0.2$. Right: This shows the SGWB characteristic strain obtained by integrating sources within the comoving volume according to Eq. (27), along with the data from PTAs. The solid curves of different colors represent the characteristic strain obtained by averaging over contributions from initial eccentricities for various dark matter model parameters. The orange dashed curve shows the best-fit case between our model and the data, with $v_{t}=396.9\,\mathrm{km\,s^{-1}}$, $\sigma_{0}/m_{\chi}=2.82\,\mathrm{cm^{2}\,g^{-1}}$, and $\psi_{0}=-2.31$. The gray dashed line represents the characteristic strain obtained without dynamical friction, assuming purely circular orbits.
@@
In Fig. 2, we plot the evolution of eccentricity as a function of the orbital semi‐major axis at an initial orbital frequency of $F_{0}=10^{-13}\,\mathrm{Hz}$, comparing the evolution predicted by the classical Chandrasekhar formula and the modified formula, Eq. (17), which incorporates additional drag effects from dark matter particles moving faster than the orbiting objects. As we can see that for small initial eccentricities (e.g. $e_{0}\lesssim 0.35$), both dynamical friction and gravitational radiation tend to circularize the orbit. In contrast, for larger initial eccentricities, the orbit is driven by dark matter friction first toward values close to unity before gravitational radiation dominates and circularizes the orbit.

### v2

@@ Abstract
The nature of dark matter is still mysterious despite various astronomical evidence. As a possible candidate, self-interacting dark matter (SIDM) can potentially resolve some issues appearing in the cold dark matter paradigm. Here we investigate how SIDM around supermassive black holes (SMBHs) in galaxy centers may form a density spike and imprint in the spectrum shape of stochastic gravitational-wave background from SMBH binaries (SMBHBs). Employing a refined dynamical friction formula and consistently evolving the orbital dynamics, we demonstrate that current pulsar timing array (PTA) data are sensitive to the cross section of SIDM with $\sigma(v)/m_{\chi}\lesssim 0.66\,\mathrm{cm}^{2}/\mathrm{g}$, comparable to other astrophysical probes. We also highlight the importance of including the eccentricity of SMBHBs in the parameter inference, which would affect the results significantly. Our findings reveal the promising potential of PTA observations in probing the nature of dark matter.
@@ I Introduction
Dark matter is one of the main components of the Universe, with its existence firmly established through multiple astronomical observations including galaxy rotation curves [1, 2], Bullet Cluster [3, 4], cosmic microwave background anisotropies [5], and large-scale structure [6]. Although at the galactic scale and beyond these observations are consistent with the cold dark matter (CDM) paradigm, at small scales or near galactic centers, the particle nature of dark matter remains uncertain. Among numerous dark matter candidates, self-interacting dark matter (SIDM) is a popular and attractive model that may behave differently at small scales [7, 8, 9, 10]. While no definitive signatures have been detected in terrestrial experiments, any new proposals to probe the nature of dark matter would be warranted.
Since the first detection of gravitational waves (GWs) [11], the advent of gravitational-wave astronomy has opened new detection windows for dark matter. Studies of GW signals with both ground‑based and space‑based detectors—such as LVK [12, 13, 14], LISA [15], and Taiji [16]—enable discrimination among a variety of dark‑matter candidates, such as CDM [17, 18], SIDM [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], and ultralight dark matter (ULDM) [32, 33, 34, 35, 36, 37, 38]. GW detection has demonstrated remarkable potential for fundamental physics [39, 40, 41, 42, 43, 26]. Notably, the search for dark matter is expected to become increasingly intertwined with gravitational-wave astronomy.
@@
While the canonical picture predicts that the characteristic strain $h_{c}$ should scale as $f^{-2/3}$ for a population of inspiraling SMBHBs [50, 51], environmental effects around black holes may alter the merger process, causing deviations of the stochastic gravitational-wave background from the ideal power-law spectrum. The NANOGrav study suggest that environmental effects may play a significant role in supermassive black hole mergers [52]. For instance, surrounding matter components such as gas, stars, and dark matter can accelerate the energy dissipation of binary systems. More studies have investigated the effects of three-body interactions [53], CDM [54, 55], ULDM [56, 57, 58, 59] and eccentric orbits on the stochastic gravitational-wave background at low frequencies [60].
@@
In this work, we adopt the SIDM framework with a velocity‑dependent scattering cross section and incorporate a refined dynamical friction formulation that captures additional drag from high‑velocity dark matter particles. We then investigate how dynamical friction affects the eccentric orbits of SMBHBs, which not only augments gravitational radiation but also induces a characteristic suppression in the SGWB’s spectrum at low frequencies relative to circular binaries. Our analysis furnishes a more realistic physical description. Through Bayesian inference applied to observed characteristic‑strain data, we obtain the parameter fit range of the SIDM model and quantify how initial orbital eccentricity reshapes the posterior distributions of model parameters.
This paper is organized as follows. Section. II presents the methodology for constructing dark matter density and velocity profiles around binary SMBHs, and we compare density profiles under different parameter configurations. Section. III details numerical computations of binary SMBH evolution incorporating eccentric orbits and dynamical friction, with particular emphasis on quantifying contributions from additional drag terms. Section. IV demonstrates how variations in SIDM parameters and initial orbital eccentricity affect the characteristic strain spectrum of stochastic gravitational waves, comparing these predictions with PTA observations. In Section. V we employ Bayesian inference and $\chi^{2}$ analysis to quantify parameter compatibility with observational data, ultimately showcasing the sensitive parameter space of SIDM models.
@@ II.1 OUTER HALO DENSITY
We consider spherically symmetric dark matter halos and use the Navarro-Frenk-White (NFW) profile [65] to characterize dark matter distribution in the outer region as a function of radial distance $r$ from the Galaxy Center,
@@
where $_{\mathrm{bulge}}, MM_{\star}$, and $M_{200}$ represent the bulge mass, stellar mass, and halo mass of the host galaxy respectively. The $M_{\mathrm{tot}}$$M_{\rm bulge}$ relation reflects the coevolution of supermassive black holes and their host galaxies [66],
@@
where $M_{\mathrm{ref}}$ is defined as $10^{12}\,h^{-1}M_{\odot}$, and the remaining parameters $C_{c},\gamma_{c},M_{c}$ are expressed as functions of redshift. We have compiled the polynomials mentioned here, along with those referenced in Eq. (6), in Appendix. A. Once we obtain $r_{s}$, we can inversely solve for $\rho_{s}$ through Eq. (7).
@@ II.2 DENSITY AND VELOCITY DISTRIBUTION OF SIDM
Here $v_{t}$ denotes the transition velocity, defined as the mass ratio between mediator and dark matter $v_{t}\equiv m_{A^{\prime}}/m_{\chi}$ and $v$ denotes the velocity dispersion of the dark matter particles. Here we do not specify the particle physics model of dark matter [71, 72, 73, 74, 75, 76, 20, 75, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88], as the relevant physical parameters are represented by $\sigma_{0}$ and $v_{t}$. It can be seen that the cross section behaves like a constant at low velocities, while at high velocities it transitions to a behavior similar to Coulomb scattering, $\sigma(v)\propto 1/v^{4}$.
Owing to the large cross section of SIDM, the exchange of momentum and energy among particles within a SIDM halo occurs rapidly. In the region with radial coordinate $r$ smaller than some value $r_{1}$, the dark matter density tends to flatten, forming an isothermal core. The boundary $r_{1}$ of the isothermal core is defined through the average interaction cross section [19, 63],
@@
where $t_{\mathrm{age}}$ represents the age of the galaxy core. This indicates that dark matter particles within the radius $r_{1}$ have at least one interaction during the galaxy merger process. We use $t_{\mathrm{age}}=1\,\mathrm{Gyr}$ in this paper for convenience. This does not affect our final analysis of the parameter $\sigma_{0}/m_{\chi}$, as we can easily rescale to different $t_{\mathrm{age}}$ by adjusting the size of $\sigma_{0}/m_{\chi}$ and compute the cross section for different characteristic velocities.
@@
where $v_{0}$ represents the velocity dispersion in the isothermal core.
@@
Figure 1: We take the parameters of the binary supermassive black hole system as the example: $M_{\mathrm{tot}}=10^{10}\,M_{\odot}$, $z=0$, $q=0.5$. Upper: the dark matter density profile is plotted for different transition velocities while fixing the dark matter self-interacting cross section $\sigma_{0}/m_{\chi}=1\,\mathrm{cm}^{2}/\mathrm{g}$. Lower: the dark matter density profile with a different cross section while fixing the transition velocity $v_{t}=500\,\mathrm{km}/\mathrm{s}$. Here, the region $r<10^{2}\,\mathrm{pc}$ corresponds to the spike, the flat middle region is the isothermal core, and the $r>10^{6}\,\mathrm{pc}$ segment belongs to the NFW region.
@@
This also represents the radius of the dark matter spike. For CDM, the density distribution of the dark matter spike can be described by a power law [91, 92, 44]. In the case of SIDM, under the nonrelativistic approximation, the density and velocity distribution can also be expressed in the form of a power law [93, 94, 63],
@@
The exponent $a$ represents the dependence of the interaction cross section on the velocity power. For the dark photon model, the interaction cross section approximates to a constant $(a=0)$ at low velocities, while at high velocities, $a=4$.
@@
We note that for certain masses of central black holes and redshifts, only a sufficiently large interaction cross section $\sigma_{0}/m_{\chi}$ can ensure that the radius of the dark matter spike $r_{\mathrm{sp}}$ exceeds the radius $r_{1}$ of the isothermal core. For dark matter spikes that have lost their isothermal core, the stable structure may be disrupted due to perturbations caused by the loss of orbital energy during binary inspiral. Meanwhile, a very large cross section may also lead to core collapse or a weakening of dynamical friction effects, causing the binary merger timescale to exceed the Hubble time [63, 95, 62].
@@
We define the transition radius $r_{t}$ such that $v(r_{t})=v_{t}$; the density distribution within the spike is given by
$\rho_{\mathrm{sp}}=\begin{cases}\rho_{0}\left(\frac{r_{\mathrm{sp}}}{r_{t}}\right)^{\frac{3}{4}}\left(\frac{r_{t}}{r}\right)^{\frac{7}{4}},&\text{if }r<r_{t}\\ \rho_{0}\left(\frac{r_{\mathrm{sp}}}{r}\right)^{\frac{3}{4}}&\text{if }r>r_{t}.\end{cases}$ (16)
@@
In Fig. 1, we construct the complete density profile and examine how the density profile responds to varying transition velocities $v_{t}$. A distinct flattened density region, corresponding to the isothermal core, appears between approximately $10^{2}$ and $10^{4}\,\mathrm{pc}$. For $v_{t}=300\,\mathrm{km}\,\mathrm{s}^{-1}$, the one–dimensional velocity dispersion is $v_{0}\simeq 373\,\mathrm{km}\,\mathrm{s}^{-1}$, so $v_{0}>v_{t}$. The corresponding transition radius $r_{t}$ lies outside the spike radius $r_{\mathrm{sp}}$, where the density slope is $\gamma=7/4$. Beyond $\sim 10^{6}\,\mathrm{pc}$, the profile reverts to the NFW form. In the lower panel of Fig. 1, with $v_{t}=500\,\mathrm{km}\,\mathrm{s}^{-1}$ fixed, we can see that increasing the cross section would broaden the flat core, which boosts particle velocities and thus lowers the density. For the same fractional variations, $v_{t}$ affects the density more strongly than $\sigma_{0}/m_{\chi}$: an order-of-magnitude increase in $v_{t}$ changes the density by roughly 3 orders of magnitude, whereas an order-of-magnitude change in $\sigma_{0}/m_{\chi}$ alters the density by only about the same order of magnitude.
@@
To verify the consistency between our model and observational results, we also conduct a benchmark analysis using the Milky Way system. By cross validating our theoretical predictions with observational constraints derived from monitoring the $S2$ stellar orbit around the Galactic Center supermassive black hole $\mathrm{Sgr}\mathrm{A}^{*}$, we find that the enclosed mass within the $S2$ orbital pericenter, even with the parameters $v_{t}=100\,\mathrm{km}/\mathrm{s},\,\sigma_{0}/m_{\chi}=0.3\,\mathrm{cm}^{2}/\mathrm{g}$ that yield the highest spike density, the density at $r=0.01\,\mathrm{pc}$ is approximately $10^{7}\,M_{\odot}/\mathrm{pc}^{3}$, and the enclosed mass does not exceed $100\,M_{\odot}$, 1 order of magnitude smaller than the empirically derived upper limit of $1200\,M_{\odot}$ [96].
@@ III DYNAMICAL FRICTION AND ORBITAL EVOLUTION
where $M_{\mathrm{bh}}$ represents the black hole mass, $v_{\mathrm{bh}}$ denotes the black hole velocity and $\Lambda$ is the Coulomb logarithm, which we take as $\log\Lambda=3$ in this work [99]. $N_{1}$ and $N_{2}$ represent the fraction of dark matter particles with velocities less than the black hole’s velocity and that with velocities between the black hole’s velocity and the escape velocity, respectively.
@@
Figure 2: This figure shows the $e(t)$ versus $a(t)$ during the dynamical friction phase of binary supermassive black holes with parameters $M_{\mathrm{tot}}=10^{10}\,M_{\odot}$, $q=0.6$, $z=2$, $v_{t}=300\,\mathrm{km}/\mathrm{s}$, and $\sigma_{0}/m_{\chi}=6\,\mathrm{cm}^{2}/\mathrm{g}$. The dashed curves represent the evolution according to the classical Chandrasekhar formula ($N_{2}=0$), while the solid curves have included the correction term $N_{2}$. The differently colored curves in the plot correspond to different initial eccentricities.
Treating all orbits as circular would undoubtedly neglect the modifications in gravitational radiation induced by eccentric orbits, thereby leading to an overestimation of the contribution from dynamical friction. In this paper, we consider orbital evolution with nonzero eccentricity, as opposed to the simplified circular orbit scenario.
@@
We use the phase-averaging method to approximate the change rates of thr $i$-th object’s energy and angular momentum in the binary system due to dynamical friction,
@@
The evolution of the orbital semimajor axis and eccentricity as functions of time are determined by the variation of the total energy and angular momentum. Detailed derivations are provided in Appendix B.
@@
We select supermassive black hole merger events within a mass range of $M_{\mathrm{tot}}/\,M_{\odot}\in[10^{6},10^{11}]$, assuming that the inspiral process of the eccentric orbit begins at an average orbital frequency $F_{0}=10^{-13}\,\mathrm{Hz}$ [101, 102]. High-resolution cosmological $N$-body simulations (IllustrisTNG100-1) [103] indicate that approximately $95\%$ of galactic merger orbits exhibit high eccentricities [102, 104], with binary systems maintaining eccentricities distributed between 0.2 and 0.8 even as they evolve into the PTA frequency band. Therefore, we consider a binary evolution system with different initial eccentricity distribution. Using numerical methods, we can determine the values of the semimajor axis $a(t)$ and eccentricity $e(t)$ at any subsequent time.
@@
Figure 3: Left: we present the variation of characteristic strain with dark matter model parameters using a single source as an example. The source parameters are $M_{\mathrm{tot},0}=10^{10}\,M_{\odot}$, $z_{0}=2$, $q_{0}=0.6$, with the event rate given by $\mathrm{d}^{3}N/(\mathrm{d}M_{\mathrm{tot}}\mathrm{d}z\mathrm{d}q)=\delta(M_{\mathrm{tot}}-M_{\mathrm{tot},0})\delta(z-z_{0})\delta(q-q_{0})$. In the figure, curves of different colors represent characteristic strain with different transition velocities and cross sections, where dashed ones correspond to an initial eccentricity of $0.8$, while the solid ones represent an initial eccentricity of $0.2$. Right: this shows the SGWB characteristic strain obtained by integrating sources within the comoving volume according to Eq. (27), along with the data from PTAs. The solid curves of different colors represent the characteristic strain obtained by averaging over contributions from initial eccentricities for various dark matter model parameters. The orange dashed curve shows the best-fit case between our model and the data, with $v_{t}=396.9\,\mathrm{km\,s^{-1}}$, $\sigma_{0}/m_{\chi}=2.82\,\mathrm{cm^{2}\,g^{-1}}$, except for the gray curve, all spectra are normalized with $\psi_{0}=-2.31$.
The gray dashed line represents the characteristic strain obtained without dynamical friction, assuming purely circular orbits.
@@
In Fig. 2, we plot the evolution of eccentricity as a function of the orbital semi‐major axis at an initial orbital frequency of $F_{0}=10^{-13}\,\mathrm{Hz}$, comparing the evolution predicted by the classical Chandrasekhar formula and the modified formula, Eq. (17), which incorporates additional drag effects from dark matter particles moving faster than the orbiting objects. As we can see, for small initial eccentricities (e.g., $e_{0}\lesssim 0.35$), both dynamical friction and gravitational radiation tend to circularize the orbit. In contrast, for larger initial eccentricities, the orbit is driven by dark matter friction first toward values close to unity before gravitational radiation dominates and circularizes the orbit.